OpenStudy (anonymous):

A bottle filled with juice weighs 8 pounds. One morning George drank 2/3 of the juice in the filled bottle. The bottle and the remaining juice weighed 3 pounds. What is the weight of the empty bottle, in pounds?

OpenStudy (texaschic101):

he drank 2/3 of the juice leaving 1/3 x = weight of the bottle 1/3 + x = 3 3/3 + x = 8 hold on....it is not equaling...let me get someone

OpenStudy (texaschic101):

@AccessDenied ...can you help.....this problem is messing with my brain

OpenStudy (texaschic101):

@johnweldon1993 ...can you help...I don't even know if I am setting this up correctly

OpenStudy (texaschic101):

@jim_thompson5910 ...can you help

OpenStudy (texaschic101):

should I have done 2/3 of 8 for what he drank ?

OpenStudy (texaschic101):

no...that doesn't make sense

jimthompson5910 (jim_thompson5910):

let x = weight of the bottle (without any juice) and y = weight of juice alone (all of the juice) we know that the bottle + juice originally is 8 lbs, so x+y = 8 solve for y to get y = 8 - x we also know that "George drank 2/3 of the juice in the filled bottle", so there's 1/3 of the juice left and this adds to the bottles weight (x) to get 3, so, x + (1/3)y = 3 x + (1/3)(8 - x) = 3 ... plug in y = 8 - x x + 8/3 - x/3 = 3 3x + 8 - x = 9 ... multiply every term by the LCD 3 to clear out the fractions ... ... ... x = ???

OpenStudy (texaschic101):

3x + 8 - x = 9 2x = 1 x = 1/2 is it 1/2 a lb ?

jimthompson5910 (jim_thompson5910):

yep, the bottle itself without any juice is 1/2 lb

OpenStudy (texaschic101):

oh...I was way off

jimthompson5910 (jim_thompson5910):

the juice alone is 8 - 1/2 = 15/2 = 7 1/2 = 7.5 lbs

jimthompson5910 (jim_thompson5910):

check: x+y = 8 0.5 + 7.5 = 8 8 = 8 works ------------------ x + (1/3)*y = 3 0.5 + (1/3)*7.5 = 3 0.5 + 2.5 = 3 3 = 3 works

OpenStudy (texaschic101):

I see what you did....it makes more sense then what I was trying to do....you have saved me AGAIN. thank you

OpenStudy (anonymous):

Thank you so much. I was stuck on it for a while.

jimthompson5910 (jim_thompson5910):

you're welcome, glad I could help

OpenStudy (texaschic101):

he is good

OpenStudy (anonymous):

thank you for calling him and for your effort :)

OpenStudy (texaschic101):

lol...I usually end up calling him when I am in a bind because I know he can help.

jimthompson5910 (jim_thompson5910):

you did good texaschic101, you just needed that extra variable

OpenStudy (texaschic101):

I always end up leaving something out :/

jimthompson5910 (jim_thompson5910):

that's ok, that's what practice is for

OpenStudy (anonymous):

can you guys help me with one more question? I posted it up but no one answered it :(

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

thank you! let me close this